OpenStudy (anonymous):

If f(x)=3x^2-6x+12, then find the minimum value of f.

OpenStudy (anonymous):

Is this multiple choice?

OpenStudy (phi):

f(x)=3x^2-6x+12 that is the equation of a parabola with the shape \(\cup\) and its minimum is at its vertex. the x value of the vertex is at \[ x = \frac{-b}{2a} \] where a , b and c are found by matching your parabola equation to \[ y = a x^2 + bx +c \]

OpenStudy (anonymous):

What does it mean when it says "then find the minimum value of f."

OpenStudy (phi):

the f(x) (or "y" if we are plotting the curve on x-y graph) is smallest at the vertex. they want the f(x) value. to find it, first find the x value of the vertex (see above) then use that number in your parabola formula to find f(x) at the vertex (which will be the min that you want)

OpenStudy (phi):

the first step is find a and b (see up above) can you do that ?

OpenStudy (anonymous):

x=1?

OpenStudy (phi):

a= 3 , b= -6 -b/(2a) is -(-6)/(2*3) = 6/6 = 1 yes x=1 now use x=1 in the formula to find the f(x) value

OpenStudy (anonymous):

so f(x)=9?

OpenStudy (phi):

I guess we should write it f(1)= 9 (which means at x=1, the function has the value 9) anyway, 9 is the smallest this parabola gets.

OpenStudy (anonymous):

Oh so that is the final answer?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank You!

OpenStudy (phi):

yw